# Some mathematical issues regarding a new approach towards quantum foundation

### v3

@@ Abstract
In this article, the weakest possible theorem giving a foundation behind the Hilbert space formalism of quantum theory is stated. The necessary postulates are formulated, and the mathematics is spelled out in detail. It is argued that, from this approach, a general epistemic interpretation of quantum mechanics is natural. Some applications to the Bell experiment and to decision theory are briefly discussed. The article represents the conclusion of a series of articles and books on quantum foundations.
@@ 1 Introduction
In a number of articles, the newest ones being Helland (2024a,b,c), this author has proposed a completely new foundation of quantum theory, a foundation based upon theoretical variables, which in a given context may be attached to an observer or to a group of communicating observers. These variables can be accessible or inaccessible. An example of a (maximal) accessible variable may be the positions $\bm{q}=(\bm{q}_{1},...,\bm{q}_{n})$ of $n$ independent particles; another example may be their momenta $\bm{p}=(\bm{p}_{1},...,\bm{p}_{n})$. A typical inaccessible variable may be the vector $(\bm{q},\bm{p})$.
From a mathematical point of view, the notion of a theoretical variable will be considered as primitive. The only requirement is that if $\theta$ is a theoretical variable and $\xi$ is a function of $\theta$, then $\xi$ is a theoretical variable. Similar, the notion of an accessible variable is primitive. Again, it is only required that if $\theta$ is accessible and $\xi$ is a function of $\theta$, then $\xi$ is accessible.
Given these primitive notions, the theory may be implemented in different ways. In this article, I will concentrate on the case where the variables are physical variables like positions, momenta or spin components. A physical variable is said to be accessible if it in principle can be measured with arbitrary accuracy.
@@
By assuming the existence of two different maximal accessible variables - in Niels Bohr’s terminology two complementary variables - and making some additional assumptions, it is shown in the above articles that essentially the whole Hilbert space apparatus results. The purpose of the present article is to look closer at the additional assumptions. It turns out that these can be considerably weakened.
What we do have to assume, is that the two variables, called $\theta$ and $\eta$ in the general theory, may be seen as functions of some basic inaccessible variable $\phi$, and that groups act on both $\phi$ and $\theta$.
@@
I will concentrate on the derivation of the Hilbert space apparatus and related derivations in this article. Assumptions that lead to the Born rule for probabilities are discussed, and the derivation is proved, in Helland (2024b). The assumptions are also discussed in Section 3 below. Derivation of the Schrödinger equation from a few postulates has been given for instance by Klein (2010).
@@
The plan of this article is as follows: In Section 2 the theory is outlined in its weakest possible version. In Section 5 some consequences, consequences to an understanding of the Bell theorem and to a new theory of decisions, are briefly discussed. In Section 6 some final remarks are given.
@@ 2 The main theorems
In simple physical examples, such a $\phi$ can easily be constructed. As a general statement covering all possible situations, Postulate 1 can also be given a religious interpretation, see Helland (2022d, 2023d). This will be further discussed elsewhere.
@@
This is a partial ordering among all theoretical variables, and also a partial ordering among the accessible ones. Note that $\phi$ from Postulate 1 is an upper bound in the accessible case. (Again, one must make precise what is meant by a function. For instance, one can let $\Omega_{\phi}$ and $\Omega_{\theta}$ be topological spaces, and concentrate on functions that are Borel-measurable.) We will say that $\theta$ is a maximal accessible variable if it is maximal with respect to this partial ordering. By Zorn’s Lemma, which is equivalent to the Axiom of Choice, maximal variables always exist. For those who do not believe in Zorn’s Lemma, we add an additional postulate.
@@
Finally, in this article, I will assume for two complementary variables $\theta$ and $\eta$::
@@
(In terms of category theory, this means that $\Omega_{\theta}$ and $\Omega_{\eta}$ are objects, and that there is a morphism from $\Omega_{\theta}$ to $\Omega_{\eta}$, and another morphism from $\Omega_{\eta}$ to $\Omega_{\theta}$. More intuitively, it means that there is a bijective function connecting $\Omega_{\theta}$ and $\Omega_{\eta}$.)
This is all we need for our first results, a Proposition and a basic Theorem.
@@
The group $M$ acting upon $\Omega_{\phi}$ need not be transitive. The points $\phi_{1}$ and $\phi_{2}$ either lie on the same orbit of $M$ or on different orbits. In the first case, there exists a $k\in M$ such that $\phi_{2}=k\phi_{1}$, so that $\eta(\phi_{1})=f(k\phi_{1})$. In the second case, let $a$ be a function of $\phi$ which characterizes the orbits. Then there exists a $\phi_{3}$ on the orbit containing $\phi_{1}$ such that $\phi_{2}=a(\phi_{3})$, and by definition $\phi_{3}=k\phi_{1}$ for some $k\in M$. In this case we get $\eta(\phi_{1})=f(\phi_{2})=f(a(\phi_{3}))=f(a(k\phi_{1}))$. Define $\xi(\phi_{1})=f(k\phi_{1})$. Then $\eta(\phi_{1})=f(a(f^{-1}(\xi(\phi_{1}))))$, which means that $\eta$ is a bijective function of $\xi$. The inverse $f^{-1}$ is well defined, since, by $\xi(\phi_{1})=f(k\phi_{1})$, and letting $\phi_{1}$ vary, the range of $\xi$ has the same category as the range of $\eta$.
@@
Theorem 1 follows now from Theorem 4 of Helland (2024a) and Proposition 2. $\Box$
This result, together with the other results of Helland (2024a,b,c) now gives a very simple alternative foundation of quantum theory. Note that this is a purely mathematical theory, and it can be interpreted in different directions. In an ordinary physical setting, it is natural to interpret the accessible theoretical variables as ordinary physical variables, but also connected to the mind of an observer or to the joint minds of a communicating group of observers. This gives a general epistemic interpretation of quantum theory, an interpretation which has QBism as a special case. Quantum theory is then a theory of an observer’s or a group of observers’ knowledge about the real world, not a theory directly about the real world.
@@
- The eigenvectors can, in the maximal case, be interpreted in terms of a question together with an answer. Specifically, it means in a context with several variables, a chosen variable $\theta$ may be associated with a question ‘What is the value of $\theta$?’ or ‘What will $\theta$ be if we measure it?’, and a specific eigenvector of $A^{\theta}$, corresponding to the eigenvalue $u$ may be identified with the answer ‘$\theta=u$.
@@
In Theorem 1 it was concluded that the relevant operators were symmetric. This is a simple property: $\langle u|Av\rangle=\langle Au|v\rangle$ for all $|u\rangle,|v\rangle$ in the domain of $A$. To use the spectral theorem in general, we need operators corresponding to the two maximal accessible variables $\theta$ and $\eta$ are self-adjoint. Then the spectral theorem is valid (Hall, 2013), and it can be used to define operators corresponding to other accessible variables.
@@
and the last two intergals are finite when $|u\rangle\in D$ and $|v\rangle\in D$. Without loss of generality assume $\||u\rangle\|=\||v\rangle\|=1$. $\Box$
@@
Since $\theta$ is maximal, all the eigenvalues are different, so $A^{\theta}$ determines uniquely the set of eigenvectors $\{\bm{v}_{j}\}$ up to phase factors. Similarly, $\{u_{i}\}$ is uniquely determined by $A^{\eta}$. The two sets of eigenvectors satisfy $\sum_{j}\bm{v}_{j}\bm{v}_{j}^{\dagger}=I$ and $\sum_{i}\bm{u}_{i}\bm{u}_{i}^{\dagger}=I$. These two sets of eigenvectors cannot be identical, for in this case $\theta$, taking the values $\lambda_{j}$, and $\eta$, taking the values $\mu_{i}$ would be bijective functions of each other. But when at least one $\bm{v}_{j}$ differs from the set of vectors $\{\bm{u}_{i}\}$, it follows from the formulae above that $A^{\theta}$ and $A^{\eta}$ do not commute. $\Box$
@@ 3 Quantum probabilities
The generalized likelihood principle from statistics holds.
The basis for nearly all statistical inference is a statistical model, a model for the data $z$ in terms of the total parameter $\theta$. This is expressed by a probability function $p(z|\theta)$, in the discrete case a point probability, and in the continuous case a probablity density. The likelihood of the data is defined as $L(\theta|z)=p(z|\theta)$, the probability function seen as a function of the parameter.
The following principle can be taken as the basis for very much statsitical inference, and in my theory it is also used as a motivation behind Born’s formula.
The generalized likelihood principle
@@
From these two postulates, assuming perfect experiments in the sense that no experimental noise is assumed, and using a version of Gleason’s theorem given by Busch (2003), the following versions of Born’s formula is proved.
Consider two experiments with maximal accessible discrete parameters $\theta^{a}$ and $\theta^{b}$, having the same number of values. By a corollary to Theorem 1 (Helland, 2024a), there are operators $A^{a}$ and $A^{b}$ in a Hilbert space $\mathcal{H}$ associated with these two parameters, the eigenvalues of the operators are the possible values of the parameters, and the corresponding normed eigenvectors are simple. Let $|a;k\rangle$ correspond to the event $\theta^{a}=u_{k}$ and $|b;j\rangle$ correspond to the event $\theta^{b}=v_{j}$.
@@
Thus, the probabilities are given by the squared norms of the probablity amplitudes $c=\langle a;k|b:j\rangle$.
@@
A referee has asked the following question: Why did we get precisely the complex Hilbert space formalism, and not the generalized probability theory, in the spirit of Khrennikov et al. (2025)? The generalized probability theory is described in detail in Barrett (2007) and Müller (2021), but we do not need the definitions here.
The answer to why quantum probablities in my theory are given by the Born formula, and not by any more generalized formula, is given by the proofs of the results above, and in particuar by a consequence of the generalized likelihood principle. In the discrete case, one can define the likelihood effects
@@
Thus, my approach towards probabilities lead to traditional quantum probabilities.
@@ 4 The epistemic interpretation and Ozawa’s intersubjectivity theorem
In other words, under the assumptions of space separabiliy and a weak reproducability condition, the two obsevers are forced to get equal measurements.
Khrennikov (2024a) recently argued that the QBism’s fundamental statement that ‘the measurement of an observable is personal’ is in contraversion with Ozawa’s theorem. His starting point is the following citation from Fuchs and Schack (2014): ‘The fundamental primitive of QBism is the concept of experience. According to QBism, quantum measurement is a theory that any agent can use to eveluate her expectations for the content of their experience.’
Following Khrennikov, Ozawi’s theorem implies that, for accurate local observations, measurement’s outcome is intersubjective which is a strong objection to QBism. There is nothing concernng personal experiences when different observers by necessity obtain the same measurement result.
This should be contrasted to my own interpretation, which can be seen as a consequence of the mathemathical theory sketched in Sections 2 and 3 above. This mathematical theory can, in a physical context be interpreted in two ways: First, the mathematics, including the Born formula, can be seen as what is given by the observations of a single observer. It should be emphasized that the probabilities in Born’s formula can be seen as probabilities calculated by the hypothetical superior actor $D$. Next, the mathematics can be interpreted as the joint observations as seen by a group of communicating observers.
Note that, in Ozawa’s theorem, the two observers are specelike separated, and do not communicate about their observations. However, spacelike separated observers may later meet and share their experiences. Then it is perfectly allowable that they find that their marurement values are equal.
In conclusion, their seems to be no contraversion between my general epistemic theory and Ozawa’s intersubjectivity theorem.
It is interesting that Khrennikov (2024b) argues that Rovelli’s well-knows relational quantum mechanics can be specified in a way where there is no contrdiction to Ozawa’s theorem. Relational quantum mechanics states that every statement about a system should be related to another system. This ‘other system’ may well be an observer or a communicating group of observers, which provides a link to my interpretation.
@@ 5 A mathematical consequence and an interpretation
The assumption that the variables are connected to the same context is crucial. In my interpretation of quantum theory, I connect the variables to the mind of an observer or to the joint minds of a communicating group of observers. Then we can consider maximal observations done in some fixed context, which also means some fixed time.
@@
In Helland (2022b, 2023b) this conclusion is applied to the observer Charlie, which observes the results of Alice and Bob in the famous Bell experiment. It is concluded that from this statement it is possible to understand that, in practice, noting that Charlie can be any observer, the violation of the CHSH inequality in practice can be understood. Note that my conclusion here is not directly a consequence of quantum mechanics, but of a series of mathematical theorems, building upon the above 4 postulates.
Another application is to decision theory. Let $O$ be faced with deciding among a finite number of actions $a_{1},...,a_{r}$. Define the decision variable $\theta$ as equal to $j$ if the action $a_{j}$ is chosen ($j=1,...,r$). Say that $\theta$ is accessible, and that the decision is accessible, if the decision can be carried out by $O$. The variable $\theta$ is called maximally accessible if the decision just can be carried out. Note that $\theta$ is finite-valued, so the theory of this article applies.
@@
Note that here, $O$ can be any person. According to my theory, we all have this limitation in our minds. The conclusion can also be generalized to the decisions made by a communicating group of individuals.

### v4

@@ Abstract
In this article, the weakest possible theorem providing a foundation for the Hilbert space formalism of quantum theory is stated. The necessary postulates are formulated, and the mathematics is spelt out in detail. It is argued that, from this approach, a general epistemic interpretation of quantum mechanics is natural. Some applications to the Bell experiment and to decision theory are briefly discussed. The article represents the conclusion of a series of articles and books on quantum foundations.
@@ 1 Introduction
In a number of articles, the newest ones being Helland (2024a,b,c), this author has proposed a completely new foundation for quantum theory, a foundation based upon theoretical variables, which in a given context may be attached to an observer or to a group of communicating observers. These variables can be accessible or inaccessible. An example of a (maximal) accessible variable may be the positions $\bm{q}=(\bm{q}_{1},...,\bm{q}_{n})$ of $n$ independent particles; another example may be their momenta $\bm{p}=(\bm{p}_{1},...,\bm{p}_{n})$. A typical inaccessible variable may be the vector $(\bm{q},\bm{p})$.
From a mathematical point of view, the notion of a theoretical variable will be considered primitive. The only requirement is that if $\theta$ is a theoretical variable and $\xi$ is a function of $\theta$, then $\xi$ is a theoretical variable. Similarly, the notion of an accessible variable is primitive. Again, it is only required that if $\theta$ is accessible and $\xi$ is a function of $\theta$, then $\xi$ is accessible.
Given these primitive notions, the theory may be implemented in different ways. In this article, I will concentrate on the case where the variables are physical variables like positions, momenta or spin components. A physical variable is said to be accessible if it can in principle be measured with arbitrary accuracy.
@@
By assuming the existence of two different maximal accessible variables, in Niels Bohr’s terminology two complementary variables, and making some additional assumptions, it is shown in the above articles that essentially the whole Hilbert space apparatus results. The purpose of the present article is to look closer at the additional assumptions. It turns out that these can be considerably weakened.
What we do have to assume is that the two variables, called $\theta$ and $\eta$ in the general theory, may be seen as functions of some basic inaccessible variable $\phi$, and that groups act on both $\phi$ and $\theta$.
@@
I will concentrate on the derivation of the Hilbert space apparatus and related derivations in this article. Assumptions that lead to the Born rule for probabilities are discussed, and the derivation is proved, in Helland (2024b). The assumptions are also discussed in Section 3 below. A derivation of the Schrödinger equation from a few postulates has been given for instance by Klein (2010).
@@
The plan of this article is as follows: in Section 2 the theory is outlined in its weakest possible version. In Section 5 some consequences, consequences for an understanding of the Bell theorem and for a new theory of decisions, are briefly discussed. In Section 6 some final remarks are given.
@@ 2 The main theorems
In simple physical examples, such a $\phi$ can easily be constructed. As a general statement covering all possible situations, Postulate 1 can also be given a religious interpretation; see Helland (2022d, 2023d). This will be further discussed elsewhere.
@@
This is a partial ordering among all theoretical variables and also a partial ordering among the accessible ones. Note that $\phi$ from Postulate 1 is an upper bound in the accessible case. (Again, one must make precise what is meant by a function. For instance, one can let $\Omega_{\phi}$ and $\Omega_{\theta}$ be topological spaces and concentrate on functions that are Borel-measurable.) We will say that $\theta$ is a maximal accessible variable if it is maximal with respect to this partial ordering. By Zorn’s Lemma, which is equivalent to the Axiom of Choice, maximal variables always exist. For those who do not believe in Zorn’s Lemma, we add an additional postulate.
@@
Finally, in this article, I will assume for two complementary variables $\theta$ and $\eta$:
@@
(In terms of category theory, this means that $\Omega_{\theta}$ and $\Omega_{\eta}$ are objects, and that there is a morphism from $\Omega_{\theta}$ to $\Omega_{\eta}$ and another morphism from $\Omega_{\eta}$ to $\Omega_{\theta}$. More intuitively, it means that there is a bijective function connecting $\Omega_{\theta}$ and $\Omega_{\eta}$.)
This is all we need for our first results: a Proposition and a basic Theorem.
@@
The group $M$ acting upon $\Omega_{\phi}$ need not be transitive. The points $\phi_{1}$ and $\phi_{2}$ either lie on the same orbit of $M$ or on different orbits. In the first case, there exists a $k\in M$ such that $\phi_{2}=k\phi_{1}$, so that $\eta(\phi_{1})=f(k\phi_{1})$. In the second case, let $a$ be a function of $\phi$ which characterises the orbits. Then there exists a $\phi_{3}$ on the orbit containing $\phi_{1}$ such that $\phi_{2}=a(\phi_{3})$, and by definition $\phi_{3}=k\phi_{1}$ for some $k\in M$. In this case we get $\eta(\phi_{1})=f(\phi_{2})=f(a(\phi_{3}))=f(a(k\phi_{1}))$. Define $\xi(\phi_{1})=f(k\phi_{1})$. Then $\eta(\phi_{1})=f(a(f^{-1}(\xi(\phi_{1}))))$, which means that $\eta$ is a bijective function of $\xi$. The inverse $f^{-1}$ is well defined, since, by $\xi(\phi_{1})=f(k\phi_{1})$, and letting $\phi_{1}$ vary, the range of $\xi$ has the same category as the range of $\eta$.
@@
Theorem 1 now follows from Theorem 4 of Helland (2024a) and Proposition 2. $\Box$
This result, together with the other results of Helland (2024a,b,c), now gives a very simple alternative foundation of quantum theory. Note that this is a purely mathematical theory, and it can be interpreted in different directions. In an ordinary physical setting, it is natural to interpret the accessible theoretical variables as ordinary physical variables, but also as connected to the mind of an observer or to the joint minds of a communicating group of observers. This gives a general epistemic interpretation of quantum theory, an interpretation which has QBism as a special case. Quantum theory is then a theory of an observer’s or a group of observers’ knowledge about the real world, not a theory directly about the real world.
@@
- The eigenvectors can, in the maximal case, be interpreted in terms of a question together with an answer. Specifically, it means in a context with several variables, a chosen variable $\theta$ may be associated with a question ‘What is the value of $\theta$?’ or ‘What will $\theta$ be if we measure it?’, and a specific eigenvector of $A^{\theta}$, corresponding to the eigenvalue $u$, may be identified with the answer ‘$\theta=u$’.
@@
In Theorem 1 it was concluded that the relevant operators were symmetric. This is a simple property: $\langle u|Av\rangle=\langle Au|v\rangle$ for all $|u\rangle,|v\rangle$ in the domain of $A$. To use the spectral theorem in general, we need operators corresponding to the two maximal accessible variables $\theta$ and $\eta$ to be self-adjoint. Then the spectral theorem is valid (Hall, 2013), and it can be used to define operators corresponding to other accessible variables.
@@
and the last two integrals are finite when $|u\rangle\in D$ and $|v\rangle\in D$. Without loss of generality assume $\||u\rangle\|=\||v\rangle\|=1$. $\Box$
@@
Since $\theta$ is maximal, all the eigenvalues are different, so $A^{\theta}$ uniquely determines the set of eigenvectors $\{\bm{v}_{j}\}$ up to phase factors. Similarly, $\{u_{i}\}$ is uniquely determined by $A^{\eta}$. The two sets of eigenvectors satisfy $\sum_{j}\bm{v}_{j}\bm{v}_{j}^{\dagger}=I$ and $\sum_{i}\bm{u}_{i}\bm{u}_{i}^{\dagger}=I$. These two sets of eigenvectors cannot be identical, for in this case $\theta$, taking the values $\lambda_{j}$, and $\eta$, taking the values $\mu_{i}$ would be bijective functions of each other. But when at least one $\bm{v}_{j}$ differs from the set of vectors $\{\bm{u}_{i}\}$, it follows from the formulae above that $A^{\theta}$ and $A^{\eta}$ do not commute. $\Box$
@@ 3 Quantum probabilities
The generalised likelihood principle from statistics holds.
The basis for nearly all statistical inference is a statistical model, a model for the data $z$ in terms of the total parameter $\theta$. This is expressed by a probability function $p(z|\theta)$, in the discrete case a point probability, and in the continuous case a probability density. The likelihood of the data is defined as $L(\theta|z)=p(z|\theta)$, the probability function seen as a function of the parameter.
The following principle can be taken as the basis for very much statisitical inference, and in my theory it is also used as a motivation behind Born’s formula.
The generalised likelihood principle
@@
From these two postulates, assuming perfect experiments in the sense that no experimental noise is assumed, and using a version of Gleason’s theorem given by Busch (2003), the following version of Born’s formula is proved.
Consider two experiments with maximal accessible discrete parameters $\theta^{a}$ and $\theta^{b}$, having the same number of values. By a corollary to Theorem 1 (Helland, 2024a), there are operators $A^{a}$ and $A^{b}$ in a Hilbert space $\mathcal{H}$ associated with these two parameters; the eigenvalues of the operators are the possible values of the parameters, and the corresponding normed eigenvectors are simple. Let $|a;k\rangle$ correspond to the event $\theta^{a}=u_{k}$ and $|b;j\rangle$ correspond to the event $\theta^{b}=v_{j}$.
@@
Thus, the probabilities are given by the squared norms of the probability amplitudes $c=\langle a;k|b:j\rangle$.
@@
A referee has asked the following question: Why did we get precisely the complex Hilbert space formalism, and not the generalised probability theory, in the spirit of Khrennikov et al. (2025)? The generalised probability theory is described in detail in Barrett (2007) and Müller (2021), but we do not need the definitions here.
The answer to why quantum probabilities in my theory are given by the Born formula, and not by any more generalised formula, is given by the proofs of the results above, and in particular by a consequence of the generalised likelihood principle. In the discrete case, one can define the likelihood effects
@@
Thus, my approach towards probabilities leads to traditional quantum probabilities.
@@ 4 The epistemic interpretation and Ozawa’s intersubjectivity theorem
In other words, under the assumptions of space separability and a weak reproducibility condition, the two observers are forced to get equal measurements.
Khrennikov (2024a) recently argued that the QBism’s fundamental statement that ‘the measurement of an observable is personal’ is in contravention of Ozawa’s theorem. His starting point is the following citation from Fuchs and Schack (2014): ‘The fundamental primitive of QBism is the concept of experience. According to QBism, quantum measurement is a theory that any agent can use to evaluate her expectations for the content of their experience.’
Following Khrennikov, Ozawa’s theorem implies that, for accurate local observations, the measurement’s outcome is intersubjective which is a strong objection to QBism. There is nothing concerning personal experiences when different observers by necessity obtain the same measurement result.
This should be contrasted with my own interpretation, which can be seen as a consequence of the mathematical theory sketched in Sections 2 and 3 above. This mathematical theory can, in a physical context be interpreted in two ways: First, the mathematics, including the Born formula, can be seen as what is given by the observations of a single observer. It should be emphasised that the probabilities in Born’s formula can be seen as probabilities calculated by the hypothetical superior actor $D$. Next, the mathematics can be interpreted as the joint observations as seen by a group of communicating observers.
Note that, in Ozawa’s theorem, the two observers are spacelike separated and do not communicate about their observations. However, spacelike separated observers may later meet and share their experiences. Then it is perfectly allowable that they find that their measurement values are equal.
In conclusion, there seems to be no contravention between my general epistemic theory and Ozawa’s intersubjectivity theorem.
It is interesting that Khrennikov (2024b) argues that Rovelli’s well-known relational quantum mechanics can be specified in a way where there is no contradiction to Ozawa’s theorem. Relational quantum mechanics states that every statement about a system should be related to another system. This ‘other system’ may well be an observer or a communicating group of observers, which provides a link to my interpretation.
It is important that my basis is a purely mathematical theory, and this theory can be interpreted in different directions. I stated above that it in a natural way leads to a general epistemic interpretation, connected to the knowledge of an observer or to the joint knowledge of a communicating group of observers. However, in some contexts, this group of communicating observers can be imagined to consist of all relevant persons in the world. In these contexts, we might as well consider an ontological interpretation of the actual aspects of quantum mechanics.
@@ 5 A mathematical consequence and an interpretation
The assumption that the variables are connected to the same context is crucial. In my interpretation of quantum theory, I connect the variables to the mind of an observer or to the joint minds of a communicating group of observers. Then we can consider maximal observations made in some fixed context, which also means some fixed time.
@@
In Helland (2022b, 2023b) this conclusion is applied to the observer Charlie, who observes the results of Alice and Bob in the famous Bell experiment. It is concluded that from this statement it is possible to understand that, in practice, noting that Charlie can be any observer, the violation of the CHSH inequality can be understood. Note that my conclusion here is not directly a consequence of quantum mechanics, but of a series of mathematical theorems, building upon the above four postulates.
Another application is to decision theory. Let $O$ be faced with deciding among a finite number of actions $a_{1},...,a_{r}$. Define the decision variable $\theta$ as equal to $j$ if the action $a_{j}$ is chosen ($j=1,...,r$). Say that $\theta$ is accessible, and that the decision is accessible, if the decision can be carried out by $O$. The variable $\theta$ is called maximally accessible if the decision can just be carried out. Note that $\theta$ is finite-valued, so the theory of this article applies.
@@
Note that here, $O$ can be any person. According to my theory, we all have this limitation in our minds. The conclusion can also be generalised to the decisions made by a communicating group of individuals.